OpenStudy (anonymous):

The minute hand of a clock is 1.4 centimeters long. How far does the tip of the minute hand travel in 40 minutes?

ganeshie8 (ganeshie8):

you may use below relation to figure out length of travel : \(l = r.\theta \) \(r = 1.4\ cm\) \(\theta = \frac{40}{60} * 2\pi\)